OpenStudy (anonymous):

-3x^2y(xy^2-2y-1) .....With steps please :)

OpenStudy (anonymous):

1min.

OpenStudy (anonymous):

I agree (1) or (2) are INSUFFICIENT. but consider both of them together (1) 2x-2y = 1 (2) x/y >1 Rephrasing (1) x-y = 1/2 x = y + 1/2 Now substituting value of x in (2) (y+1/2)/y >1 Solving: 1+ 1/2y > 1 so, 1/2y > 0 it means y >0 since x = y+1/2 s0 X >0 as well. Hope it helps

OpenStudy (anonymous):

That has nothing to do with the question she asked?

OpenStudy (anonymous):

You need to use the distributive property to simplify this. Take the term outside the parentheses and multiply it by each term inside.